# Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning

### v1

@@ 6 Conclusion & Limitation
Finding that the theoretical justification of maximizing $I(Z;M)$ can lead to performance improvement has been lacking, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, after scrutinizing this optimization framework, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound, proving that by reining the task representation shift, it is possible to achieve monotonic performance improvements. To make it practical, we propose a tuning-free algorithm RETRO with only adding one line of code compared to the backbone. Extensive experiments have validated the superiority of RETRO on training stability, asymptotic performance and training-time consumption. Though we choose classifier-based loss to train the task encoder, the idea that reins the task representation shift is general, which is shown in Section 5.3. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift, which is also stated in Section 5.3.
@@ 7.2 Missing Proofs
$=\mathbb{E}_{m,x}[|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|$
$+|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|]$ (21)
@@
Corollary 7.5 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (35)
@@
Proof: Let $Z_{1}$ has already been trained on $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ to get the monotonic performance improvement guarantee.
By applying Lemma 7.2, the $L_{1}$ deviation of the empirical distribution $Z_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{|\mathcal{\mathbb{Z}}|}-2)\exp{(-(N+k)\epsilon^{2}/2)}$ (36)
@@
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{N+k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (37)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (38)
@@ 7.3 Implementation Details
In this section, we report the implementation details of RETRO and all the baselines.
  • FOCAL[li2020focal]. We adopt the open-source code of FOCAL [^1] https://github.com/FOCAL-ICLR/FOCAL-ICLR/. To better show the performance of FOCAL in each environment, we tune the parameter $\beta$ in distance metric learning from $[0.1,0.5,1]$ and choose the best result in our baseline comparison.
  • CSRO[gao2024context]. We adopt the open-source code of CSRO [^2] https://github.com/MoreanP/CSRO. To better show the performance of CSRO in each environment, we tune the club loss weight and task representation loss weight from $[10,20,30]\times[5,10,15]$ and choose the best result in our baseline comparison.
  • UNICORN[li2024towards]. We obtain the source code of UNICORN from the authors. To better show the performance of UNICORN in each environment, we tune the parameter $\frac{\alpha\beta-1}{(1-\alpha\beta)}$ from $[0.1,0.3,0.7,1,1.5]$ and choose the best result in our baseline comparison.
  • Classifier. Built upon the code base of FOCAL, Classifier only replaces the distance metric learning loss with cross-entropy loss, achieving tuning-free property.
  • RETRO. Built upon the code base of Classifier, RETRO only adjusts the update frequency of the task encoder to half of the origin, by adding merely one row of code.
@@
Configurations Point-Robot Humanoid-Dir Reach Ant-Dir Button-Press Dial-Turn
dataset size 2100 9e4 3e5 9e4 3e5 3e5
task representation dimension 5 64 5 5 5 5

### v2

@@ 7 Conclusion & Limitation
Finding that the performance improvement guarantee of the COMRL training framework remains underexplored, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that the previous optimization framework can be linked with the general RL objective of maximizing the expected return. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation. Based on this finding, we propose a new issue called task representation shift and prove that it is possible to achieve monotonic performance improvements with proper context encoder updates. With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift.
@@ 8.1 Useful Lemmas
Proof:
$\begin{split}|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|&=|\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}(p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a))r(s,a)|\\ &\leq R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}|p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a)|\\ &=2R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\\ \end{split}$ (14)
According to Theorem 2 (Return Bound) in [zhang2023how],
$\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\leq(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (15)
We bring this result back and can draw the conclusion.
@@
Proof:
According to the inequality for $L_{1}$ deviation of the empirical distribution [weissman2003inequalities], they conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\phi(\pi_{P})\epsilon^{2}/4)$, where $\phi(p)=\frac{1}{1-2p}\log\frac{1-p}{p}$ and $\pi_{P}=\max_{A\in\mathcal{A}}\min(P(A),1-P(A))$. The paper [weissman2003inequalities] point out that $0\leq\pi_{P}\leq 1/2,\forall P$. Based on this, we can derive the function $\phi(p)$ and find that $\phi(p)$ is monotonically decreasing on $(0,1/2)$. Therefore, we can get $\phi(p)\geq\phi(1/2)=2$. Then, we bring this result back and can conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\epsilon^{2}/2)$.
@@ 8.2 Missing Proofs
$=\mathbb{E}_{m,x}[|V^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}|+|V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{m}|]$ (22)
@@
Theorem 8.5 (Monotonic performance improvement guarantee on training process).
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (36)
@@
Proof:
Let $Z_{1}$ have already been trained by maximizing $I(Z;M)$ to some extent. Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ samples from the training dataset by maximizing $I(Z;M)$ to get the monotonic performance improvement guarantee.
We begin with the following inequality:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|+|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ (37)
where $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
According to Assumption 4.9, we have:
$|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (38)
Now, we move our focus on solving $|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $Z^{empirical}_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (39)
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (40)
@@
$|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (41)
Hence, we can get:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (42)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ (43)
$k\geq\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (44)
Since we need the least number of samples that update the context encoder, we take $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$. ∎
Corollary 8.6 (Monotonic performance improvement condition for pre-training scheme).
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (45)
Proof: Denote $J^{pretrain}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the policy. Similarly, denote $J^{pretrain}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the policy.
$J^{pretrain}(\theta_{2})-J^{pretrain}(\theta_{1})$ (46)
$=J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (47)
$J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (48)
$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))]+\epsilon^{mutual}_{12}$ (49)
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1}))]$ (50)
$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}]$ (51)
To get the monotonic performance improvement, if we consider the condition more strictly, we need the term within the expectation of Eq. (51) to be larger than 0. Hence, we can get:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{1}{2\kappa L_{z}}\epsilon^{mutual}_{12}=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (52)
@@ 8.3 Implementation Details
In this section, we report the implementation details of all the objectives in our main paper.
  • Contrastive-based [li2020focal]. We adopt the open-source code of FOCAL [^3] https://github.com/FOCAL-ICLR/FOCAL-ICLR/.
  • Reconstruction-based [li2024towards]. We obtain the source code of UNICORN from the authors and adopt UNICORN-0 as the reconstruction-based objective training framework.
  • Cross-entropy-based. Built upon the code base of FOCAL, cross-entropy-based only replaces the distance metric learning loss with cross-entropy loss.
@@
Configurations Reach Ant-Dir Button-Press Dial-Turn
dataset size 3e5 9e4 3e5 3e5
task representation dimension 5 5 5 5